OpenStudy (anonymous):

can someone take me through this equation and why the answer is (none): 3r+1/r+3 +2 = 5r-2/r+3

OpenStudy (anonymous):

\[\frac{ 3r+1 }{ r+3 } + 2 = \frac{ 5r-2 }{ r+3 }\]

OpenStudy (lgbasallote):

first...multiply ALL terms by r+3 to get rid of the denominator what do you get?

OpenStudy (anonymous):

5r+7=5r-2 > 5r-5r=-2+7 is 5r-5r =r?

OpenStudy (lgbasallote):

5r - 5r is 0

OpenStudy (anonymous):

so whenever r or x or whatever they are looking for are 0 then there is no answer for the equation?

OpenStudy (lgbasallote):

yup

OpenStudy (anonymous):

thank you very much!

OpenStudy (lgbasallote):

welcome